OpenStudy (anonymous):

(2x³y²)-² (-̅x̅̅̅²̅y̅^̅5̅) ̅̅̅

OpenStudy (anonymous):

@nikato

OpenStudy (ranga):

= 2^(-2) * x^(-6) * y^(-4) / (-x^2 * y^5) = 1 / (2^2 * x^6 * y^4 * -x^2 * y^5) = -1 / (4 * x^8 * y^9) = -1/(4x^8y^9)